OpenStudy (anonymous):

11x + 3y = 75 2x + y = 15 Using the two equations above, solve for x.

OpenStudy (anonymous):

Either substitute or eliminate, solve for one of the variables, and then plug it into the other equation to solve for the other. In this case, you could do that y = 15-2x, and then plug that into 11x+3y = 75. Or, you could eliminate by doing\[(11x-6x) + (3y-3y) = (75-15)\], and then just solve for x.